A sample of Argon has a volume of 3.2 L and the pressure is 0.9 atm. If the final temperature is 34 c, the final volume is 1 L, and the final pressure is 6 atm, what was the initial temperature of the argon in Kelvin?


1
Expert's answer
2021-02-24T03:40:10-0500

Combined gas law:

P1V1/T1=P2V2/T2 ;


P1=0.9 atm

V1=3.2 L

T1 -- ?

P2=6 atm

V2=1 L

T2=34 oC= 307 K


T1=P1V1T2/P2V2=0.9*3.2*307/(6*1)=147.36 K


Answer: the initial temperature of the argon was 147.36 K